HELLO, please Find the area of a triangle with sides 6.4 km, and 8.9 km and included angle 63° ?
Answer by rothauserc(4718)

Find the area of a triangle with sides 6.4 km, and 8.9 km and included angle 63° ?
The solution requires the use of the Side Angle Side (SAS) formula for the area of a triangle:
SAS = (1/2)*(side1 * side2)*sine(angle)
SAS = (1/2)*(6.4 * 8.9) * sine(63)
SAS = (1/2)*(56.96)*(.89)
SAS = 28.48 * .89 = 25.3 square km